OpenStudy (anonymous):

- 2pi * (x/2pi) = 5pi how to put this in a math equation

OpenStudy (anonymous):

\[2\pi*\frac{ x }{ 2\pi }=5\pi\]

OpenStudy (anonymous):

did i do it right?

OpenStudy (anonymous):

yes